1.  Calculate the force between a 0.05 C charge and a 0.03 C charge 2 meters apart.
1
Expert's answer
2021-05-14T09:41:25-0400

"F=k\frac{q_1q_2}{x^2}=9\cdot10^9\cdot\frac{0.05\cdot0.03}{2^2}=3 375 000\ (N)" . Answer